# Energy Dynamics Powered by Traction and Stress Control Formation and Motion of +1/2 Topological Defects in Epithelial Cell Monolayers

### v1

@@ Abstract
In confluent cell monolayers, patterns of cell forces and motion are systematically altered near topological defects in cell shape. In turn, defects have been proposed to alter cell density, extrusion, and invasion, but it remains unclear how the defects form and how they affect cell forces and motion. Here, we studied +1/2 defects, and, in contrast to prior studies, we observed both tail-to-head and head-to-tail defect motion occurring at the same time in the same cell monolayer. We quantified the cell velocities, the tractions at the cell-substrate interface, and stresses within the cell layer near +1/2 defects. Results revealed that both traction and stress are sources of activity within the epithelial cell monolayer, with their competition defining whether the cells inject or dissipate energy and determining the direction of motion of +1/2 defects. Interestingly, patterns of motion, traction, stress, and energy injection near +1/2 defects existed before defect formation, suggesting that defects form as a result of spatially coordinated patterns in cell forces and motion. These findings reverse the current picture, from one in which defects define the cell forces and motion to one in which coordinated patterns of cell forces and motion cause defects to form and move.
@@ Introduction
In embryonic morphogenesis, collective cell migration, and cell extrusion, cells often align the long axes of their bodies with those of their neighbors, producing nematic order. 1, 2, 3, 4 Flaws in the order, called topological defects, can affect the cell layer in important ways, such as by altering the local cell density, inducing formation of holes, causing extrusion of cells from the cell layer, and altering rates of cancer cell invasion.5, 2, 6, 7 Consistent with these phenomena, patterns of cell velocities are systematically altered near the defects.8, 5, 2, 9, 4, 10 Additionally, near topological defects, the stress state within the monolayer is altered, with clear local gradients in intercellular stresses present.2, 10, 11
@@
Given the systematic patterns of cell velocity, density, and stresses occurring at defects, topological defects provide a useful means of studying the relationships between cell shape, velocity, and force production. However, it is not yet clear how the defects in nematic order form and how exactly they are related to the cell forces and the collective migration. It is commonly assumed that cell monolayers are well described by the theory of active nematic materials, in which the amount of nematic order results from a competition between the tendency for cells to align with their neighbors and the magnitude of active stresses produced by the cells.12, 13, 14 In this description, defects form as the active stresses increase relative to the tendency for alignment between neighboring cells. Theories often assume that the force generating machinery within each cell aligns with the major axis of the cell body.5, 2, 6, 10 Following this assumption, one can predict cell velocity patterns near defects, but, paradoxically, epithelial cells have been observed to move opposite to predictions.10 The cause of this discrepancy has not yet been determined, but it may be that prior studies did not fully consider the forces at the cell-cell and cell-substrate interfaces.
@@
When applied to cell monolayers, theories based on active nematic materials have typically assumed that the traction between the cell and the substrate acts as a damping friction that impedes motion,15, 16, 17, 18, 19 with some theory being supported by experimental data.5, 6 For epithelial cells, it is not obvious that cell-substrate traction would act as a passive friction, especially given prior observations that tractions can have strong effects on cell motion and aspect ratio. 20 Additionally, some experiments have observed situations in which tractions align with, rather than against, the cell velocity.21, 22, 23 As alignment of traction and velocity would inject energy into the system, the observations of alignment between traction and velocity indicates that cell-substrate tractions are another potential source of activity. This complicates understanding of what parameters control nematic order and formation of defects, as now the tendency for alignment between neighboring cells would be balanced against both the active stresses within the cell layer and the active tractions at the cell-substrate interface. Hence, further investigation is required to uncover the relationships between nematic defects, the stresses, the tractions, and the collective cell motion.
@@
Here, we studied the forces leading to formation and motion of +1/2 topological defects in islands of Madin Darby canine kidney (MDCK) epithelial cells. We quantified both the stresses within the cell layer and the tractions at the cell-substrate interface. We observed some defects moved in the tail-to-head direction, consistent with prior observations for this cell type.10 Interestingly, other defects in the same monolayer moved in the opposite direction, from the head to the tail. Through detailed analysis of the forces and kinematics in these experiments, we found that the tail-to-head defects moved against gradients in intercellular stress and with the direction of traction. This observation indicates that in addition to gradients in stress, traction is also a major controller of +1/2 topological defect motion in epithelial cell layers. To combine stresses, tractions, and flows into a single parameter, we quantified the power density, with results showing that cells near tail-to-head defects injected energy into the system, whereas cells near head-to-tail defects predominantly dissipated energy. Finally, we show that patterns of motion, stresses, tractions, and energy injection associated with +1/2 defects existed in the cell layer before the defects formed, which suggests that formation and motion of nematic defects may result from patterns in force and motion occurring before the defects form.
@@ Results
Figure 1: Topological defects in MDCK cell layers. (a) Phase contrast image of a confluent MDCK cell island. (b) Cell orientation plot with yellow lines indicating local orientation. +1/2 and -1/2 defects are indicated by the comet tails and tripods, respectively. Scale bar is 250 µm. (c, d) Representative +1/2 defects in MDCK cell layers moving in the tail-to-head direction (c) and head-to-tail direction (d). Yellow lines indicate cell orientation, white arrows indicate cell velocity, and color maps indicate vorticity. Note that the vorticity has opposite signs on either sides of the defect tails. (e) Local coordinate system used to analyze all +1/2 defects having the origin $(x=0,y=0)$ located at the defect head and the positive $x$ direction pointing along the defect tail. Arrows show motion of the defect in the tail-to-head (blue) and head-to-tail (red) directions. (f) Fraction of defects with tail-to-head and head-to-tail motion in cell layers under the different treatments. (g) Distribution of average $x$ component of cell velocity ($<\!v_{x}\!>$) for 111 identified +1/2 defects. For each defect, $v_{x}$ is averaged over cells located in the tail region ($1000<x<$ 100 µm, $y=0$). Out of 111 +1/2 defects, there were 58 head-to-tail and 53 tail-to-head moving defects.
@@
We seeded MDCK cells in islands of 1 mm diameter on polyacrylamide substrates of Young’s modulus 6 kPa, as in our prior work.22, 20 Cells were seeded to confluence and imaged at a relatively low density of approximately 2400 cells/mm2, which results in a larger aspect ratio compared to higher densities. 20 In some experiments, we modulated the cell forces using CN03 (2 µg/mL), which activates Rho thereby increasing actomyosin contraction, and, separately, using CN02 (40 ng/ml), which activates Rac/Cdc42 activity thereby perturbing organization of actin within the cytoskeleton. Both treatments increased the average traction and the cell contractile stresses but did not alter the average cell migration speed (Supplemental Fig. S1). Following prior efforts, 2, 5, 6 we imaged the cell layers by optical microscopy over time and identified topological defects. From an image of the cell layer (Fig. 1a), the orientation field was measured, and topological defects were identified by singularity points. 5, 10, 2, 7, 24 The topological defect charge, defined as the rotation experienced by the orientation field along a path encircling the defect, was either +1/2 or -1/2, represented by comet and tripod symbols, respectively (Fig. 1b).
@@
As isolated -1/2 defects are not expected to move due to their three-fold symmetry, we focused on +1/2 defects for this study. Identified defects were verified by quantifying the corresponding cell velocity and vorticity fields using image correlation 25 and comparing to theoretical predictions, namely the +1/2 defects are predicted to have a pair of vortices on either side of the defect’s tail. For +1/2 defects, the average cell velocity can be either in the tail-to-head or head-to-tail direction, with the direction thought to depend on whether the system is, respectively, extensile, wherein the cells create pressure and push along their long axis, or contractile, wherein the cells produce tension and pull along their long axis. Near the +1/2 defects, the cells followed the predicted patterns of vorticity and velocity, and, interestingly, both tail-to-head and head-to-tail cell velocities were observed in all cell islands (Fig. 1c,d, Supplemental Fig. S2). After quantifying cell velocities, we tracked the motion of the +1/2 defects themselves. The direction of motion of the defect (tail-to-head or head-to-tail) matched that of the cell velocity (Supplemental Fig. S2 and Videos 1-3). Prior studies 2, 9, 10, 24 interpreted the direction of defect motion as indicating extensile or contractile behavior. The use of defect velocity to distinguish between extensile and contractile has been recently drawn into question, 26 however, so hereafter we use the language “tail-to-head” and “head-to-tail” to describe cell migration and defect motion.
[…]
@@
We analyzed the strain rate near the defects, focusing on the strain rate along the tail of the defect, $\dot{\varepsilon}_{xx}$. We first determined the time point at which each defect formed (which we define as $t=0$). Next, we averaged $\dot{\varepsilon}_{xx}$ for all head-to-tail and all tail-to-head moving defects at the time point immediately after defect formation ($t=0$) and 1 hr before defect formation ($t=-1$ hr). At $t=0$, the strain rates were consistent with prior reports,10 namely, at the defect (near the origin), the strain rate was positive for head-to-tail moving defects and negative for tail-to-head defects (Fig. 2b, f). The difference in sign indicates that the cells elongate (for head-to-tail defects) or shorten (for tail-to-head) along the axis of each defect. Magnitudes of strain rate were on the order of 10% hr-1. Remarkably, for both types of +1/2 defects, the patterns of strain rate were also present at $t=-1$ hr, before defect formation (Fig. 2a, e, and Supplemental Fig. S3).
@@
Figure 2: Strain rate and stress fields near +1/2 defects. (a–d) For head-to-tail moving defects, the strain rate $\dot{\varepsilon}_{xx}$ and the normal stress $\sigma_{xx}$ were averaged over all 58 defects and plotted at $t=-1$ hr (1 hr before defect formation) and $t=0$ hr (the first time point after defect formation). (e–h) For tail-to-head moving defects, $\dot{\varepsilon}_{xx}$ and $\sigma_{xx}$ were averaged over 53 defects and shown for $t=-1$ hr and $t=0$. For all plots, the coordinate system is the same as in Fig. 1e.
@@
Next, we examined the stress state within the cell layer near +1/2 defects, measuring the in-plane stresses using monolayer stress microscopy.27, 28, 29 At the scale of the entire cell monolayer, there was a slight tendency for alignment between the orientations of the cells and of first principal stress, though the alignment was not perfect (Supplemental Fig. S4), as reported recently.26 Near +1/2 defects, past literature reported a pattern wherein the stress increased along the axis of the defect in the head-to-tail direction.2, 10 As we are focused on motion along the axis of the defect (which is the $x$ direction in our chosen coordinate system), we considered the average normal stress in that direction, i.e., $\sigma_{xx}$, and we averaged $\sigma_{xx}$ over many defects. For both types of defects, there was a positive gradient in $\sigma_{xx}$ along the direction from head to tail (Fig. 2d,h, Supplemental Fig. S5), which is consistent with prior observations.2, 10
@@
The fact that the gradient in stress is the same for both tail-to-head and head-to-tail motion is puzzling, because, given the opposite direction of motion, one would expect the gradients to be opposite as well. In particular, in the case of tail-to-head motion, the cells moved against gradients in stress, i.e., away from regions of high tension, which is akin to moving toward regions of high pressure. Although this puzzling observation has been made before,10 the underlying cause remains unclear. To investigate more deeply, we plotted $\sigma_{xx}$ an hour before defect formation, considering the same local coordinate system as of the formed defect. Similar to the strain rate, the patterns in stresses at $t=-1$ hr matched those at $t=0$ (Fig. 2c,g). These findings are distinct from prior observations in that the prior studies reported spatial patterns of strain rate and stress after the defects formed, but not before.2, 10
@@
Next, we quantified the variability between the different defects. We calculated the average $x$ component of velocity and the gradient $\partial\sigma_{xx}/\partial x$ along the tail region of each defect, averaged the data for each defect, and plotted the results in 2D heat maps. Although there was notable variability in $<\!\partial\sigma_{xx}/\partial x\!>$, for defects moving in both the head-to-tail and tail-to-head directions, $<\!\partial\sigma_{xx}/\partial x\!>$ was positive for most defects (Supplemental Fig. S6). We also considered that analyzing only $\sigma_{xx}$ is incomplete, as doing so neglects the other components of the stress tensor. Following prior studies, 2, 10 we calculated the average principal stress, $(\sigma_{1}+\sigma_{2})/2$. Results showed similar trends, with the averaged gradient of $(\sigma_{1}+\sigma_{2})/2$ being positive along the tail of the defect for both head-to-tail and tail-to-head defects (Supplemental Fig. S6). We also computed the gradient of stress more rigorously. As equilibrium is mathematically written as the divergence of stress, we computed the expression corresponding to equilibrium in the $x$ direction, $\partial\sigma_{xx}/\partial x+\partial\sigma_{xy}/\partial y$. Again, the sign was positive, for both head-to-tail and tail-to-head moving defects (Supplemental Fig. S7). These data confirmed that tail-to-head defects moved against gradients in stress, and, interestingly, patterns of both strain rate and stress existed before defect formation.
@@
Figure 3: Traction applied by the substrate to the cells near +1/2 defects. (a) Phase contrast images over time of a representative +1/2 tail-to-head moving defect. Lines indicate cell orientation. (b) Vector plots of velocity (white arrows) and traction (red arrows) at the same time points as in panel a. The arrows typically point in the same direction, indicating a tendency for alignment between velocity and traction for tail-to-head moving defects. (c, d) Plots of traction averaged over 58 head-to-tail moving defects at $t=-1$ hr and at $t=0$ hr. (e, f) Plots of traction averaged over 53 tail-to-head moving defects at $t=-1$ hr and at $t=0$ hr.
Motion against the stress gradient means that the stress impedes the cell motion, which is only possible if there exist other forces that propel the cells against the stress gradient. A potential force that can propel the cells against the stress gradient is the substrate-to-cell traction. We measured tractions applied by the substrate to the cells using traction force microscopy.30, 31, 32, 33 Most models of active nematic fluids15, 16, 17, 18, 19 assume the traction to be a viscous drag that acts as a friction. In such a case, the traction (applied by the substrate to the cell) would point exactly opposite to the direction of cell velocity. In contrast to this expectation, visual examination of a representative +1/2 tail-to-head defect showed that the fields of traction and velocity tended to align (Fig. 3a,b). The alignment was especially notable in the tail region the defect. To study this observation in multiple defects, we averaged the tractions for all head-to-tail and tail-to-head defects and plotted the results as vector plots for $t=-1$ hr and $t=0$ hr (Fig. 3c-f). For both cases, the traction always pointed toward the head of the defects. These data suggest that for tail-to-head defects, the traction acts in the same direction as the motion, meaning it is propulsive. We confirmed this observation further by plotting histograms of the angle between traction and velocity, $\theta$. For all cells across multiple monolayers, the distribution of $\theta$ was approximately uniform (Supplemental Fig. S8). A uniform distribution of $\theta$ also occurred near defect-free regions containing counter-rotating vortices (Supplemental Fig. S8). Near defects, $\theta$ showed a tendency towards 0 or 180, depending on whether the motion was tail to head or head to tail, respectively. For tail-to-head defects, the alignment between traction and velocity along the tails indicates the opposite of friction, implying that the traction propelled the cells forward (Supplemental Fig. S8). Hence, there exists an apparent competition between stress gradient and tractions, with stresses controlling the motion of head-to-tail defects and tractions controlling the motion of tail-to-head defects.

### v2

# Traction and Stress Control Formation and Motion of +1/2 Topological Defects in Epithelial Cell Monolayers

@@ Abstract
In confluent cell monolayers, patterns of cell forces and motion are systematically altered near topological defects in cell shape. In turn, defects have been proposed to alter cell density, extrusion, and invasion, but it remains unclear how the defects form and how they affect cell forces and motion. Here, we studied +1/2 defects, and, in contrast to prior studies, we observed the concurrent occurrence of both tail-to-head and head-to-tail defect motion in the same cell monolayer. We quantified the cell velocities, the tractions at the cell-substrate interface, and the stresses within the cell layer near +1/2 defects. Results revealed that both traction and stress are sources of activity and dissipation within the epithelial cell monolayer, with the direction of motion of +1/2 defects depending on whether energy is injected by stresses or tractions. Interestingly, patterns of motion, traction, stress, and energy injection near +1/2 defects existed before defect formation, suggesting that defects form as a result of spatially coordinated patterns in cell forces and motion. These findings introduce a new focus, on coordinated patterns of force and motion that lead to defect formation and motion.
Key Words
collective cell migration, topological defects, traction force microscopy, energy injection and dissipation, active matter
@@ Introduction
In embryonic morphogenesis, collective cell migration, and cell extrusion, cells often align the long axes of their bodies with those of their neighbors, producing nematic order. 1, 2, 3, 4 Flaws in the order, called topological defects, can affect the cell layer in important ways, such as by altering the local cell density, inducing formation of holes, causing extrusion of cells from the cell layer, and altering rates of cancer cell invasion.5, 2, 6, 7 Consistent with these phenomena, patterns of cell velocities are systematically altered near the defects.8, 5, 2, 9, 4, 10 Additionally the stress state within the monolayer is altered near topological defects, with the presence of clear local gradients in intercellular stresses.2, 10, 11
@@
Given the systematic patterns of cell velocity, density, and stresses occurring at defects, topological defects provide a useful means of studying the relationships between cell shape, velocity, and force production. However, it is not yet clear how the defects in nematic order form and how exactly they are related to the cell forces and the collective migration. It is commonly assumed that cell monolayers are well described by the theory of active nematic materials, in which the amount of nematic order results from a competition between the tendency for cells to align with their neighbors and the magnitude of active stresses produced by the cells.12, 13, 14 In this description, defects form as the active stresses increase relative to the tendency for alignment between neighboring cells. Theories often assume that the force generating machinery within each cell aligns with the major axis of the cell body.5, 2, 6, 10 Following this assumption, one can predict cell velocity patterns near defects, but, paradoxically, epithelial cells that produce contractile (tensile) stresses have been observed to move opposite to predictions, as if their active stresses were extensile (pushing).2, 9, 10 The cause of this discrepancy has not yet been determined, but it may be that prior studies did not fully consider the forces at the cell-cell and cell-substrate interfaces.
@@
When applied to cell monolayers, theories based on active nematic materials have typically assumed that the traction between the cell and the substrate acts as a damping friction that impedes motion,15, 16, 17, 18, 19 with some theory being supported by experimental data.5, 6, 20 For epithelial cells, it is not obvious that cell-substrate traction would act as a passive friction, especially given prior observations that tractions can have strong effects on cell motion and aspect ratio. 21 Some models for active nematics have considered that traction may have an additional part that propels the cell forward,22, 23 as have other models for collective cell migration that do not consider the cell orientation field.24, 25, 26, 27 Additionally, some experiments in epithelial cell monolayers have observed situations in which tractions align with, rather than against, the cell velocity.28, 26, 29 As alignment of traction and velocity would inject energy into the system, the observations of alignment between traction and velocity indicates that cell-substrate tractions are another potential source of activity. This complicates understanding of what parameters control nematic order and formation of defects, as now the tendency for alignment between neighboring cells would be balanced against both the active stresses within the cell layer and the active tractions at the cell-substrate interface. Hence, further investigation is required to uncover the relationships between nematic defects, the stresses, the tractions, and the collective cell motion.
@@
Here, we study the forces leading to formation and motion of +1/2 topological defects in islands of Madin Darby canine kidney (MDCK) epithelial cells. We quantify the stresses within the cell layer and the tractions at the cell-substrate interface. We observe that some defects move in the tail-to-head direction, consistent with prior observations for this cell type. 10 Interestingly, other defects in the same monolayer move in the opposite direction, from the head to the tail. Through detailed analysis of the forces and kinematics in these experiments, we find that the tail-to-head defects move against gradients in intercellular stress and with the direction of traction. This observation identifies that traction can act as a driving force to propel +1/2 defects against gradients in stress. To combine force and displacement into a single parameter, we quantify the power densities associated with stresses and traction, showing that for tail-to-head moving defects, the energy for cell motion comes primarily through the cell-substrate tractions, whereas for head-to-tail defects, the energy comes from transmission of stresses applied by neighboring cells. Finally, we show that patterns of motion, stresses, tractions, and energy injection associated with +1/2 defects existed in the cell layer before defect formation, suggesting that formation and motion of nematic defects may result from patterns in force and motion occurring before the defects form.
@@ Results
We seeded MDCK cells in islands of 1 mm diameter on polyacrylamide substrates of Young’s modulus 6 kPa, as in our prior work.26, 21 Cells were seeded to confluence and imaged at a relatively low density of approximately 2400 cells/mm2, which results in a larger aspect ratio compared to higher densities. 21 In some experiments, we modulated the cell forces using CN03 (2 µg/mL), which activates Rho thereby increasing actomyosin contraction, and, separately, using CN02 (40 ng/ml), which activates Rac/Cdc42 activity thereby perturbing organization of actin within the cytoskeleton. Both treatments increased the average traction and the cell contractile stresses but did not alter the average cell migration speed (Supplemental Fig. S1). This finding is consistent with prior studies that showed no effect of CN03 on average cell speed.21, 30 Following prior efforts, 2, 5, 6 we imaged the cell layers by optical microscopy over time and identified topological defects. From an image of the cell layer (Fig. 1a), the orientation field was measured, and topological defects were identified by singularity points. 5, 10, 2, 7, 31 The topological defect charge, defined as the rotation experienced by the orientation field along a path encircling the defect, was either +1/2 or -1/2, represented by comet and tripod symbols, respectively (Fig. 1b).
Figure 1: Topological defects in MDCK cell layers. (a) Phase contrast image of a confluent MDCK cell island. (b) Cell orientation plot with yellow lines indicating local orientation. +1/2 and -1/2 defects are indicated by the comet tails and tripods, respectively. Scale bar is 250 µm. (c, d) Representative +1/2 defects in MDCK cell layers moving in the tail-to-head direction (c) and head-to-tail direction (d). Yellow lines indicate cell orientation, white arrows indicate cell velocity, and color maps indicate vorticity. Note that the vorticity has opposite signs on either sides of the defect tails. (e) Local coordinate system used to analyze all +1/2 defects having the origin $(x=0,y=0)$ located at the defect head and the positive $x$ direction pointing along the defect tail. Arrows show motion of the defect in the tail-to-head (blue) and head-to-tail (red) directions. (f) Fraction of defects with tail-to-head and head-to-tail motion in cell layers under the different treatments. (g) Distribution of average $x$ component of cell velocity ($v_{x}\!>$) for 111 identified +1/2 defects. For each defect, $v_{x}$ is averaged over cells located in the tail region ($0<x<100$ µm, $y=0$). Out of 111 +1/2 defects, there were 58 head-to-tail and 53 tail-to-head moving defects.
@@
As isolated -1/2 defects are not expected to move due to their three-fold symmetry, we focused on +1/2 defects for this study. Identified defects were verified by quantifying the corresponding cell velocity and vorticity fields using image correlation 32 and comparing to theoretical predictions, namely the +1/2 defects are predicted to have a pair of vortices on either side of the defect’s tail. For +1/2 defects, the average cell velocity can be either in the tail-to-head or head-to-tail direction, with the direction thought to depend on whether the system is, respectively, extensile, wherein the cells create pressure and push along their long axis, or contractile, wherein the cells produce tension and pull along their long axis. Near the +1/2 defects, the cells followed the predicted patterns of vorticity and velocity, and, interestingly, both tail-to-head and head-to-tail cell velocities were observed in all cell islands (Fig. 1c,d, Supplemental Fig. S2). After quantifying cell velocities, we tracked the motion of the +1/2 defects themselves. The direction of motion of the defect (tail-to-head or head-to-tail) matched that of the cell velocity (Supplemental Fig. S2 and Videos 1-3). Prior studies 2, 9, 10, 31 interpreted the direction of defect motion as indicating extensile or contractile behavior. The use of defect velocity to distinguish between extensile and contractile has been recently drawn into question, 33 however, so hereafter we use the language “tail-to-head” and “head-to-tail” to describe cell migration and defect motion.
Prior reports with epithelial cell types indicated that the direction of defect motion was primarily tail-to-head; the motion was in the opposite direction for highly elongated cell types or epithelial cells with the cell-cell adhesion protein E-cadherin knocked out.8, 2, 5, 9, 10, 31, 11 By contrast, in our experiments, defects moved in either direction. To quantify these observations, we located +1/2 defects in 27 different islands across all treatments. We kept for further analysis only the defects that exhibited counter-rotating vortices (as in Fig. 1c,d) and were located at least 100 µm away from the island boundary. Our selection criteria resulted in a total of 111 defects, which we rotated such that the tail pointed along the positive $x$ direction of the local coordinate system, with the origin being located at the head of each defect (Fig. 1e). We then computed the average $x$ component of velocity of cells along the tail of each defect in positions $0<x<100$ µm, $y=0$, which we refer to as $v_{x}\!>$. (Hereafter, we use brackets $\cdot\!$> to indicate quantities that have been averaged along the defect tails.) Interestingly, the data showed that neither tail-to-head nor head-to-tail motion was predominant (Fig. 1f). A histogram of $v_{x}\!>$ for all defects was nearly symmetric about zero but with two well separated peaks (Fig. 1g), indicating a typical cell migration speed of a few µm/hr with motion either head to tail or tail to head.
@@
We analyzed the strain rate near the defects, focusing on the strain rate along the tail of the defect, $\dot{\varepsilon}_{xx}$. We first determined the time point at which each defect formed (which we define as $t=0$). Next, we averaged $\dot{\varepsilon}_{xx}$ for all head-to-tail and all tail-to-head moving defects at the time point immediately after defect formation ($t=0$) and 1 hr before defect formation (t=1t=-1 hr). At $t=0$, the strain rates were consistent with prior reports,10 namely, at the defect (near the origin), the strain rate was positive for head-to-tail moving defects and negative for tail-to-head defects (Fig. 2b,f). The difference in sign indicates that the cells elongate (for head-to-tail defects) or shorten (for tail-to-head) along the axis of each defect. Magnitudes of strain rate were on the order of 10% hr-1. Remarkably, for both types of +1/2 defects, the patterns of strain rate were also present at t=1t=-1 hr, before defect formation (Fig. 2a,e, and Supplemental Fig. S3).
@@
Figure 2: Strain rate and stress fields near +1/2 defects. (a–d) For head-to-tail moving defects, the strain rate $\dot{\varepsilon}_{xx}$ and the normal stress $\sigma_{xx}$ were averaged over all 58 defects and plotted at t=1t=-1 hr (1 hr before defect formation) and $t=0$ hr (the first time point after defect formation). (e–h) For tail-to-head moving defects, $\dot{\varepsilon}_{xx}$ and $\sigma_{xx}$ were averaged over 53 defects and shown for t=1t=-1 hr and $t=0$. For all plots, the coordinate system is the same as in Fig. 1e.
@@
Next, we examined the stress state within the cell layer near +1/2 defects, measuring the in-plane stresses using monolayer stress microscopy.34, 35, 36 At the scale of the entire cell monolayer, there was a slight tendency for alignment between the orientations of the cells and of first principal stress, though the alignment was not perfect (Supplemental Fig. S4), as reported recently.33 Near +1/2 defects, past literature reported a pattern wherein the stress increased along the axis of the defect in the head-to-tail direction.2, 10 As we are focused on motion along the axis of the defect (which is the $x$ direction in our chosen coordinate system), we considered the average normal stress in that direction, i.e., $\sigma_{xx}$, and we averaged $\sigma_{xx}$ over many defects. For both types of defects, there was a positive gradient in $\sigma_{xx}$ along the direction from head to tail (Fig. 2d,h, Supplemental Fig. S5), which is consistent with prior observations.2, 10
@@
The fact that the gradient in stress is the same for both tail-to-head and head-to-tail motion is puzzling, because, given the opposite direction of motion, one would expect the gradients to be opposite as well. In particular, in the case of tail-to-head motion, the cells moved against gradients in stress, i.e., away from regions of high tension, which is akin to moving toward regions of high pressure. Although this puzzling observation has been made before,10 the underlying cause remains unclear. To investigate more deeply, we plotted $\sigma_{xx}$ an hour before defect formation, considering the same local coordinate system as of the formed defect. Similar to the strain rate, the patterns in stresses at t=1t=-1 hr matched those at $t=0$ (Fig. 2c,g). These findings are distinct from prior observations in that the prior studies reported spatial patterns of strain rate and stress after the defects formed, but not before.2, 10
@@
Next, we quantified the variability between the different defects. We calculated the average $x$ component of velocity and the gradient $\partial\sigma_{xx}/\partial x$ along the tail region of each defect, averaged the data for each defect, and plotted the results in 2D heat maps. Although there was notable variability in $\partial\sigma_{xx}/\partial x\!>$, for defects moving in both the head-to-tail and tail-to-head directions, $\partial\sigma_{xx}/\partial x\!>$ was positive for most defects (Supplemental Fig. S6). We also considered that analyzing only $\sigma_{xx}$ is incomplete, as doing so neglects the other components of the stress tensor. Following prior studies, 2, 10 we calculated the average principal stress, $(\sigma_{1}+\sigma_{2})/2$. Results showed similar trends, with the averaged gradient of $(\sigma_{1}+\sigma_{2})/2$ being positive along the tail of the defect for both head-to-tail and tail-to-head defects (Supplemental Fig. S6). We also computed the gradient of stress more rigorously. As equilibrium is mathematically written as the divergence of stress, we computed the expression corresponding to equilibrium in the $x$ direction, $\partial\sigma_{xx}/\partial x+\partial\sigma_{xy}/\partial y$. Again, the sign was positive for both head-to-tail and tail-to-head moving defects (Supplemental Fig. S7). These data confirmed that tail-to-head defects moved against gradients in stress, and, interestingly, patterns of both strain rate and stress existed before defect formation.
@@
Motion against the stress gradient means that the stress impedes the cell motion, which is only possible if there exists another source of activity that can propel the cells against the stress gradient. A potential source of activity is the substrate-to-cell traction. We measured tractions applied by the substrate to the cells using traction force microscopy.37, 38, 39, 40 Most models of active nematic fluids15, 16, 17, 18, 19 assume the traction to be a viscous drag that acts as a friction. In such a case, the traction (applied by the substrate to the cell) would point exactly opposite to the direction of cell velocity. In contrast to this expectation, visual examination of a representative +1/2 tail-to-head defect showed that the fields of traction and velocity tended to align (Fig. 3a,b). The alignment was especially notable in the tail region of the defect. To study this observation in multiple defects, we averaged the tractions for all head-to-tail and tail-to-head defects and plotted the results as vector plots for t=1t=-1 hr and $t=0$ hr (Fig. 3c-f). For both cases, the traction always pointed toward the head of the defects, i.e., in the negative $x$ direction, which is consistent with the fact that the traction balances the stress gradient, which was in the positive $x$ direction (Fig. 2).
Figure 3: Traction applied by the substrate to the cells near +1/2 defects. (a) Phase contrast images over time of a representative +1/2 tail-to-head moving defect. Lines indicate cell orientation. (b) Vector plots of velocity (white arrows) and traction (red arrows) at the same time points as in panel a. The arrows typically point in the same direction, indicating a tendency for alignment between velocity and traction for tail-to-head moving defects. (c, d) Plots of traction averaged over 58 head-to-tail moving defects at t=1t=-1 hr and at $t=0$ hr. (e, f) Plots of traction averaged over 53 tail-to-head moving defects at t=1t=-1 hr and at $t=0$ hr.
These data suggest that for tail-to-head defects, the traction acts in the same direction as the motion, meaning it is propulsive. We confirmed this observation further by plotting histograms of the angle between traction and velocity, $\theta$. For all cells across multiple monolayers, the distribution of $\theta$ was approximately uniform (Supplemental Fig. S8). A uniform distribution of $\theta$ also occurred near defect-free regions containing counter-rotating vortices (Supplemental Fig. S8). Near defects, $\theta$ showed a tendency towards 0 or 180, depending on whether the motion was tail to head or head to tail, respectively. For tail-to-head defects, the alignment between traction and velocity along the tails indicates the opposite of friction, implying that the traction propelled the cells forward (Supplemental Fig. S8). In summary, for head-to-tail moving defects, stress gradients propelled the motion, and tractions acted as a resisting friction; for tail-to-head, tractions propelled, and stress gradients resisted.